From sth@ast.cam.ac.uk Thu Mar 24 18:38:43 2011 Date: Wed, 23 Mar 2011 18:41:26 +0000 From: Simon Hodgkin To: Mike Irwin Subject: minutes to hours CASU meeting Wednesday March 23rd 11:30-13:00 CASU meeting area. Present: MJI, AKY, EGS, JPE, JRL, MR, NAW, SC, STH Apologies: RGM, JCR Agenda: 1. Actions from last meeting 2. Comments on WFAU minutes 3. Hardware update 4. Meetings and telecons 5. Data archives update 6. Optical/NIR processing 7. WFCAM update 8. VISTA update 9. AOB Minutes ======= 1. Actions from the last meeting -------------------------------- AKY - has now obtained an UltraVISTA image from HJM. Investigations start here. <<<< STH - put the i-band vs 2MASS colour relations on to our web pages. <<<< (JPE commented that he has implemented them in the SADTool). AKY - has been investigating flatfield bevaiour with time and will turn this into a report with additional input from MJI. Issues for consideration are: do bad pixels change over time ? how do the detector-to-detector offsets vary ? is there any evidence for changing spatially dependent features ? how do the means, medians, noise properties, of the pixel values change. <<<< EGS, AKY, MJI - completed the v1.0-v1.1 migration for data up until the end of Feb28 and flagged it as ready to go. MJI - computed detecotr level offsets and illumination corrections for VISTA. He comented that Z and Y seem to be more variable (at the few percent level), while JHK effects seem to be pretty small. EGS, AKY - As part of the EVALSO test file transfers from ESO Garching to CASU, EGS and AKY found that a total of 8 files were missing from 13 nights of test transfer data. MJI - still needs to put the picture showing the variation in MJD across a tile on the web with some accompanying text. <<<< MJI - The only reliable way to guarantee overall tile QC is to make use of the provenance information in the headers to track the QC information in the pawprint stacks making up the tile. He said he would add a note to this effect on the relevant web pages. **Not sure if you did this or not** <<<< JRL - get better. He did. 2. Comments on WFAU Minutes --------------------------- No comments 3. Hardware Update ------------------ ** Need input from Mike ** APM44 requires a kernel module upgrade. The APM19 problem has been resolved for now. But APM19e had an excitingly different problem. We also ordered a new raid which has been interesting. The new raid is fibre-channel with ~20Tbytes. 4. Meetings and Telecons ------------------------ MJI noted that the next VISTA IoT is scheduled for April 14th. MJI reported that there was nothing much to report from a VISTA VMC Skype conference held on 24th Feb. Interestingly they didn't know about the VISTA recoating. It turns out that none of the PI's were told until recently. Various of us attended various PLATO and Euclid meetings. MJI will be attending NAM to talk about M31, and also reminded us about the Herts VVV/VPHAS/EGAPS meeting in July. 5. Data archives update ------------------------ All good. 6. Optical/NIR processing -------------------------- A few bits and bobs from Gerry Gilmore (INT WFC) to process. We also heard news that the VST test camera is on and delivering Image Quality of around 0.6 arcseconds. EGE has been asked if we will also process some VST SN survey data for them. MJI has been talking to ESO re processing all of the data. The VPHAS+ filters are still in Garching. 7. WFCAM update ---------------- The weather continued to be a little grumpy. MR noted a problem with headers for data taken on 16th+17th Feb. A tiny number (10) of Detector2 files are missing headers. This showed up initially as a WCS problem. MR is not yet sure if this is propagated to later nights (pending more processing). There is a clue - it may well be correlated with frames that have a nonsense pixel value. AY - action fix that bloody script. A full homespace is an unhappy homespace. (que?) All data are now processed to the end of Feb 2011, and has been transferred to ROE. We had some discussion over the resubmission of the extended UKIDSS survey to the UKIRT Board. The main goals are to finish the UHS and extend the UDS, with some extra GPS narrow band imaging (all available on the UKIDSS wiki). We hope this will happen. Processed data was provided for Ricky Smartt's PI parallax programme. MJI saw some results presented at the last UKIDSS meeting and was suitably impressed (and not just that the CASU pipeline centroiding is better than a bespoke PSF solution). 8. VISTA update ---------------- Ingestion and processing is all done apart from the first 3 days of March. V1.1 products are released up to and including end of January. The Feb data is processed - pending photometric chip-to-chip offsets and any spatial dependent stuff. MJI pointed out that a few of the per-directory Semaphore files were missing. We were unsure as to where ROE are with the copying. We talked about releasing the data. For the VVV we are writing catalogues to disk and sending them to Garching - then ESO will pass on the disks to them. We discussed again the network transfer of files between Garching and here - using the EVALSO thingy (from Paranal to Garching). EGS commented that the median transfer speed is 12-16 Gbytes per minute (Faster than a USB connection). One while night takes about 1-2 hours. Last month's minutes carried an eroneous number. Interestingly some disks and some people were lost in Paris. MJI asked if we ever heard back from Matt Jarvis re his narrow-band data - and specifically if he is happy. There was a lot of shaking of heads. We discussed the ESO OB status and grading parameters. V1.1 now has the correct keywords in the headers - specifically the OB overall grade is now assigned to a tile rather than the first pawprints values. We also had some discussion re. ESO PHASE3. ** need input here**. One interesting thing concerning the test file transfer is that we are limited to only 8 threads for uploads to ESO. MJI and EGE will investigate whether we can have 1 or 2 or many more. <<<< Action on NAW: find out what the bandwidth to ESO is ? Without upsetting anyone. <<<< The Deliverables are tiles and tiled catalogues for everyone. And pawprint images + catalogues for VHS. JRL was asked if heever replied to a software request to trigger on N images rather than an OB when the OB takes a long time to execute, and therefore the QC info is slow to update at the telescope. Jim says "this is not my problem", and has communicated this to WHU and explained exactly who's problem this is. Mike asked us to think about more pretty pictures for Richard Hook following a request from same. M15 and Carina were mentioned. Should probably be an action on someone I guess. <<<<< RGM has announced the formation of the VHSUCDSWG. Honestly. STH, EGS and JPE will (and did) sit (sat) in this afternoon. MJI discussed a tiny V1.1-V1.2 upgrade. There is a small binning issue for the very edges of the illumination correction tables. For the time being we are using the nearest bin at the very edges, but in the next release we will have more and better bins. JRL offered a bet: 10 quid that we still get empty disks from ESO. STH naively took him up on this. ESO - are you listening - I split it with you. JRL agreed to add photometric zeropoint and calibration comments to the fits headers following the ELIXIR (CFH) example. 9. CASUHELPs ------------- Is there any update on who is using using CASUTOOLS ? Herve Bouy mailed us for help: the VISTA pipeline doesn't build on Darwin (Mac OSX) ! AOB --- Some general flimflam about spectroscopic pipelines for the upcoming megaGaiaFLAMESUVES survey. ---------------------------------------------------------------------------------------------- Simon Hodgkin Cambridge Astronomical Survey Unit, Tel: +44 (0)1223 766657 Institute of Astronomy, Madingley Road Fax: +44 (0)1223 337523 Cambridge, CB3 0HA. UK ----------------------------------------------------------------------------------------------